# Rethinking Programmed I/O for Fast Devices, Cheap Cores, and Coherent Interconnects

### v1

@@ 4.4. Discussion
In most implementations, this assumption typically holds in order to maximize the memory bandwidth utilization and minimize request latency, and also to simplify reasoning about deadlock freedom of the coherence protocol itself. However, it may be the case that the cache stripes transactions across a limited number of independent units and which might deadlock if both A and B were mapped to the same unit. Were this to be the case, it could be avoided by careful placement of A and B in the physical address space. This is actually the case in our FPGA device implementation, described in Section 4.5.
4.5. Implementation in Enzian
@@
We implemented both device-homed protocols in the Enzian platform, which we use for the evaluation in Section 5. The implementation uses SystemVerilog and exploits an existing “Directory Controller” on Enzian’s FPGA (Ramdas, 2024) which manages the coherent links to the CPU and handles transient states in the ECI protocol.
@@
ECI provides all the messages we require to implement our messaging protocols. In particular, it includes the ability for the device to return a line to the CPU cache in Exclusive state, even though the CPU requested it in Shared. This optimization (discussed in Section 4.3) dramatically reduces latency, as shown in Section 5.1.
@@
In our case, no flow-control issues arise since ECI uses separate virtual channels for request and reply messages, removing the chance of deadlock. However, other complications arise: ThunderX-1 generates Global Synchronization (GSync) messages which function as global barriers. A GSync blocks any subsequent transactions until all those started before the GSync have retired. But the invalidates sent are not awaited by the core they are still completed allowing the protocol to progress.
@@
The ThunderX-1 divides its last-level cache functionality into a set of units called TADs, each of which can handle up to 16 simultaneous cache transactions. Two consecutive cache lines will be mapped to the different TADs by L2 to ensure that the operations can proceed independently.
@@
On this CPU, barriers are needed to reliably order reads and writes. For example, it is critical that the core’s write buffer is drained into L1 cache (which is write-through to the L2) before a subsequent load from the core signals to the FPGA that the line can be pulled. The ARM architecture’s weak memory model requires a DMB barrier instruction between the writes and the read, and on the ThunderX-1 this is sufficient to ensure ordering observed at the FPGA.
On the FPGA, control (interception) is performed based on AXI read requests, since both Shared and Exclusive operations trigger these requests. The AXI writes go directly to the application. If the read is the result of the Exclusive, then the dummy data is sent back to the CPU as here is no fresh result yet.
@@ 5. Evaluation
In this section, we compare PIO over a coherent interconnect with two other reference points: PIO directly over PCIe– essentially, the CPU copying data to and from a region of PCIe address space – and descriptor-based DMA over PCIe. All use the same Enzian hardware platform. Coherent PIO uses the Enzian native ECI, while the two PCIe techniques connect the CPU and FPGA PCIe interfaces back-to-back with a loopback cable as described in Section 3, effectively turning the Enzian FPGA into a conventional PCIe FPGA card. DMA experiments all use the Xilinx XDMA hardware IP on the FPGA.
We explore three different application scenarios: synchronous RPC to an accelerator on the FPGA, a closely-coupled high-speed NIC, and a stream processor which moves some of its dataflow operator graph to the FPGA.
5.1. RPC to an accelerator
The first experiment measures null-RPC latency when the CPU invokes a function on the device, passing arguments, and obtaining a result. The protocol shown in 4(a) and detailed in Section 4.3 is used (i.e. the cache lines used for communication are homed on the FPGA).
Figure 5. Round trip ECI RPC latency, plus software latency between two sockets
@@ 5.1. RPC to an accelerator
For comparison, we also show an implementation of the FastForward protocol (Giacomoni et al., 2008a) exchanging cache lines between sockets in a dual-socket ThunderX-1-based Gigabyte R150-T61 server with similar CPU and DDR specification to Enzian; this achieves a median latency of about 1150ns.
The optimized protocol is performing 2 round-trip message exchanges over ECI, which has a one-way latency at the link later of about 150ns. The rest of the overhead (300ns) is mostly incurred in the protocol processing in the directory controller. The FPGA is clocked at about 300MHz.
A significant penalty is incurred when the results of the RPC are returned in Shared, resulting in further round trips over ECI to (unnecessarily) invalidate the line at the FPGA before sending the next message. This suggests that future coherent interconnects would benefit greatly from allowing this kind of operation.
@@
5.2. RPC over ECI vs. PCIe PIO and DMA
We now compare the RPC latency achieved using our protocol over ECI with PIO over PCIe and the Xilinx XDMA hardware using descriptors.
{listing}
[t] {minted} [ frame=lines, framesep=2mm, baselinestretch=1.2, bgcolor=LightGray, fontsize=, linenos, ] c // map 4KiB of PCIe memory for read/write int fd = open(”/sys/bus/pci/devices/…/resource0”, O_RDWR); void *pcie_bar = mmap(NULL, 4096, PROT_READ — PROT_WRITE, MAP_SHARED, fd, 0);
// write RPC arguments to PCIe memory *(volatile rpc_args_t *)pcie_bar = rpc_args; __sync_synchronize();
// read RPC result from PCIe memory rpc_res_t rpc_res = *(volatile rpc_res_t *)pcie_bar; __sync_synchronize(); Example RPC using PIO over PCIe on Linux. Error handling not shown for brevity.
For PIO over PCIe we pre-map the relevant PCIe aperture (BAR) into user space on the CPU and measure the time to write a value to this space and read a result back, as illustrated in Section 5.2. For XDMA numbers we use Xilinx XDMA benchmark (Xilinx DMA Benchmark, 2024), which essentially performs the same operation using descriptor-based DMA. In both these cases, the FPGA is configured to map the request and response into a write to and a read from onboard Block RAM, respectively.
@@
The results are shown in Figure 6. Up to payloads of 8KiB, the latency of XMDA-based communication is completely dominated by descriptor setup and manipulation, and remains largely independent of payload size. Faster PCIe interconnects *recall that this is a PCIe Gen3 x8 link) would only further emphasize this.
Latency over ECI is dramatically lower for smaller payloads. Unsurprisingly, it is constant up to 128 bytes (the size of a cache line on Enzian) and then increases linearly, since the protocol simply replicates each cache line exchange until all the data has been transferred. It is only at 8KiB that ECI RPC latency exceeds that of XMDA. One might argue that the transfer of multiple cache lines over ECI might be improved by clever pipelining and locating the lines in physical memory so as to exploit parallelism in the directory controllers, but we do not explore that in this paper.
Figure 6. RPC latency for different payload sizes
[…]
On this platform, using common techniques to improve performance like vector instructions has limited or no benefit for PIO over PCIe. For writes, the hardware already performs coalescing, and we see transactions of up to 512 bits regardless of vector size (confirmed using a logic analyzer). For reads, the 128 bit limit remains, and is responsible for the increase in latency between 16-byte and 32-byte payloads.
@@
5.3. Network interface
[…]
We implement 3 variations of a 100Gb/s NIC on Enzian. The first (7(a)) is a conventional approach connecting one of Enzian’s 100Gb Ethernet MACs fairly directly to the XDMA engines, resulting in a NIC which is programmed using descriptor rings. The second design (7(b)) buffers packets in onboard SRAM on the FPGA and allows the CPU to read and write packet and control data using PCIe reads and writes. The final approach (7(c)) uses a custom module on the FPGA to bridge between the ECI directory controller and the Ethernet MAC, using the protocol in 4(b) to deliver network packets to the CPU in cache lines. In all cases we clock the core NIC logic at 250 MHz.
@@ 5.3. Network interface
Figure 8 shows receive latency for different packet sizes for each of the NIC implementations. As in the RPC experiment, DMA latency is fairly stable across all packet sizes, suggesting that the latency is dominated by various DMA overheads, such as descriptor setup and manipulation.
Note that system calls here actually result in very little overhead (only a few microseconds); the overhead is actually dominated by cache misses manipulating DMA data structures in main memory.
@@
Figure 9 shows transmit latency. In this case, PCIe PIO competes well with ECI across all packet sizes due to combining posted writes. PIO over ECI shows a slightly higher latency for small packets, due to it needing 2 round-trips over ECI for each 128 bytes, versus a single PCIe round trip. Both PIO solutions perform significantly better than DMA across all packet sizes.
These results show a clear advantage of using PIO over DMA, even for larger packets. The poor receive performance of PCIe-based PIO for receiving packets is largely due to a limitation of the PCIe design, that disappears for transmit.
5.4. Worked Example: Offloading Timely Dataflow
In this example we apply the techniques and results outlined so far to a practical workload: Timely Dataflow (Timely Dataflow, 2020), a real-time stream-processing system. We compare the baseline performance to a partially FPGA-offloaded implementation under three CPU–FPGA communication scenarios: DMA over PCIe, PIO over PCIe, and PIO over ECI.
@@ 5.4. Worked Example: Offloading Timely Dataflow
To test the limits, we construct a worst-case scenario: a large (31-element) graph of simple operators (filters). The large graph maximizes communication overhead in the form of progress tracking data, while the simple operators leave the FPGA with only a small performance lead over the CPU. We want to predict whether PIO can make FPGA offload work in regimes where the fixed overheads of DMA make it impractical, for example where small batch sizes are needed to ensure data freshness (as in Differential Dataflow).
@@
As shown by the $r^{2}$ values, the fit for each dataset is excellent. We can thus model the time taken to process a batch as a fixed component, or overhead, independent of the amount of data processed (e.g. running the Timely scheduler, or setting up DMA descriptors) plus a component which scales linearly with the size of the batch.
@@
Alternatively or additionally, bringing the fixed cost down reduces the smallest practical batch. If the relative overhead could be brought below 500ns/batch then with parallel cache lines giving a 1ns/B transfer cost, FPGA offload using ECI-PIO would outperform the CPU from a batch size of 128B, or a single cache line. With such an optimized protocol, ECI-PIO would outperform all other cases until a batch size of 142KiB, above which PCIe-DMA would overtake it giving a net offload benefit across all batch sizes.
Figure 10. DMA and PIO performance for Timely Dataflow FPGA offload with a graph size 31

### v2

@@ 5. PIO over a coherent interconnect
In most implementations, this assumption typically holds in order to maximize the memory bandwidth utilization and minimize request latency, and also to simplify reasoning about deadlock freedom of the coherence protocol itself. However, it may be the case that the cache stripes transactions across a limited number of independent units and which might deadlock if both A and B were mapped to the same unit. Were this to be the case, it could be avoided by careful placement of A and B in the physical address space. This is actually the case in our FPGA device implementation, described in Section 5.
Implementation in Enzian
@@
We implemented the two device-homed protocols over Enzian and evaluate them in Section 6. The implementations use SystemVerilog and SpinalHDL (contributors, 2024) and exploit an existing “Directory Controller” on Enzian’s FPGA (Ramdas, 2023) which manages the coherent links to the CPU and handles transient states in the ECI protocol.
@@
ECI provides all the messages we require to implement our messaging protocols. In particular, it includes the ability for the device to return a line to the CPU cache in Exclusive state, even though the CPU requested it in Shared. This optimization dramatically reduces latency, as shown in Section 6.1.
@@
In our case, no flow-control issues arise since ECI uses separate virtual channels for request and reply messages, removing the chance of deadlock. However, other complications arise: ThunderX-1 generates Global Synchronization (GSync) messages which function as global barriers. A GSync blocks any subsequent transactions until all those started before the GSync have retired. Fortunately, any invalidations sent are not blocked by core, allowing the protocol to progress regardless.
@@
The ThunderX-1 divides its last-level cache functionality into a set of units called TADs, each of which can handle up to 16 simultaneous cache transactions. We ensure that consecutive cache lines are mapped to different TADs to ensure that the operations can proceed independently.
@@
The ARMv8-A weak memory model means that barriers are needed to reliably order reads and writes. For example, it is critical that the core’s write buffer is drained into L1 cache (which is write-through to the L2) before a subsequent load from the core signals to the FPGA that the line can be pulled, and this requires a DMB barrier instruction between the writes and the read. On the ThunderX-1 this is sufficient to ensure ordering observed at the FPGA.
@@ 6. Evaluation
We compare PIO over a coherent interconnect with two other reference points: PIO directly over PCIe– essentially, the CPU copying data to and from a region of PCIe address space – and descriptor-based DMA over PCIe. All use the same Enzian hardware platform. Coherent PIO uses the Enzian native ECI, while the two PCIe techniques connect the CPU and FPGA PCIe interfaces back-to-back with a loopback cable as described in Section 3, effectively turning the Enzian FPGA into a conventional PCIe FPGA card. DMA experiments all use the Xilinx XDMA hardware IP on the FPGA.
We explore three different application scenarios: synchronous invocation to an accelerator on the FPGA, a closely-coupled high-speed NIC, and a stream processor which moves some of its dataflow operator graph to the FPGA.
6.1. Invoking functions on an accelerator
We first measure latency when the CPU invokes a function on the device, passing arguments, and obtaining a result. We use the protocol shown in 5(a), whereby the cache lines are homed on the FPGA.
Figure 6. Invocation latency over ECI vs. FastForward
@@ 6.1. Invoking functions on an accelerator
For comparison, we also show an implementation of the FastForward protocol (Giacomoni et al., 2008) exchanging cache lines between CPU sockets in a dual-socket ThunderX-1-based Gigabyte R150-T61 server with similar CPU and DDR specification to Enzian; this achieves a median latency of about 1150ns.
The optimized protocol is performing 2 round-trip message exchanges over ECI, which has a one-way latency at the link layer of about 150ns. The rest of the overhead (300ns) is mostly incurred in the protocol processing in the directory controller. The FPGA is clocked at about 300MHz.
A significant penalty is incurred when the results of the call are returned in Shared, resulting in further round trips over ECI to (unnecessarily) invalidate the line at the FPGA before sending the next message. This suggests that future coherent interconnects would benefit greatly from allowing the home node to return a line in Exclusive.
@@
6.2. Invocation over ECI vs. PCIe PIO and DMA
Figure 7. Invocation latency for different payload sizes
We now compare the latency achieved in the previous section with both PIO over PCIe and the Xilinx XDMA hardware using descriptors. For DMA numbers we use the Xilinx XDMA benchmark (Xilinx DMA Benchmark, 2024), which essentially performs the same operation using descriptor-based DMA. In both these cases, the FPGA is configured to map the request and response into a write to, and a read from, onboard Block RAM.
@@
The results are shown in Figure 7. Up to payloads of 8KiB, the latency of XMDA-based communication is dominated by descriptor setup and manipulation, and remains largely independent of payload size. Faster PCIe interconnects (this is a PCIe Gen3 x8 link) would amplify this effect.
Latency over ECI is dramatically lower for smaller payloads. Unsurprisingly, it is constant up to 128 bytes (the size of a cache line on Enzian) and then increases linearly, since the protocol simply replicates each cache line exchange until all the data has been transferred. It is only at 8KiB that ECI invocation latency exceeds that of XMDA. Arguably the transfer of multiple cache lines over ECI might be improved by clever pipelining and locating the lines in physical memory so as to exploit parallelism in the directory controllers, but we do not explore that in this paper.
Relative to ECI, PIO over PCIe performs poorly except for payloads under 16 bytes in size. This is due to limitations in PCIe that we have discussed in Section 4, notably the limited read bus width.
@@
This experiment is closed-loop and single-threaded, and so achieved throughput is proportional to the round-trip latency. Fully saturating the raw link bandwidth (in all cases) would require many parallel invocations at once, which would go against our goal of supporting applications where latency is paramount.
6.3. Network interface adapter
@@
We implement 3 variations of a 100 Gb/s NIC in Enzian. The first (8(a)) is a conventional approach connecting one of Enzian’s 100 Gb/s Ethernet MACs directly to the XDMA engines, resulting in a NIC programmed using descriptor rings. The second design (8(b)) buffers packets in onboard SRAM on the FPGA and allows the CPU to read and write packet and control data using PCIe reads and writes. The final approach (8(c)) uses a custom module on the FPGA to bridge between the ECI directory controller and the Ethernet MAC, using a slight variant of the protocol in 5(b) to deliver network packets to the CPU in multiple cache lines. In all cases we clock the core NIC logic at 250 MHz.
@@
Table 1. NIC implementations: latency percentiles on selected packet sizes. Higher tail latencies are marked in bold.
Size Percentiles RX ($\mu$s) Percentiles TX ($\mu$s)
P50 P95 P99 P100 P50 P95 P99 P100
PCIe DMA
64 65.39 66.36 67.65 100.01 10.06 10.35 10.59 16.49
1536 64.77 65.59 66.21 133.84 10.89 11.12 11.33 30.84
9600 65.89 67.12 68.05 123.61 15.73 16.05 16.29 41.99
PCIe PIO
64 3.25 3.26 3.27 3.39 0.34 0.36 0.38 4.80
1536 72.89 72.93 72.96 73.05 1.82 1.84 1.86 6.40
9600 450.28 450.35 450.38 451.10 9.91 10.01 10.07 10.14
ECI PIO
64 1.05 1.06 1.07 1.17 1.06 1.12 1.14 1.18
1536 7.24 7.29 7.39 7.43 3.09 3.26 3.50 3.59
9600 39.43 39.48 39.50 39.55 9.07 9.19 9.65 9.95
To minimize disruptions from the Linux scheduler, we follow the common practice to free one CPU core from kernel tasks and IRQ processing (Akkan et al., 2012), with the isolcpus function together with core pinning using taskset. In this setup, the only disturbance from Linux would be the 250 Hz timer interrupt present on all cores. While a custom, completely tickless kernel would remove this disturbance completely, it is not commonly deployed in production environments due to performance complications. We use such a tickless kernel to measure tail latency below, but other experiments use a stock generic kernel from Ubuntu.
Figure 9 shows receive latency for different packet sizes for each of the NIC implementations. As in the invocation experiment, DMA latency is fairly stable across all packet sizes, suggesting that the latency is dominated by various DMA overheads, such as descriptor setup and manipulation.
Note that system calls here actually result in little overhead (only a few $\mu$s); the overhead is dominated by cache misses manipulating DMA data structures in main memory.
In contrast, PIO over both ECI and PCIe offer much lower latency for small packet sizes up to 1024 bytes DMA, highlighting high efficiency gains in PIO compared to DMA. However, the latency of PCIe PIO quickly degrades for larger packets due to the non-posted read problem discussed in Section 4, reaching over 350 $\mu$s for 8 KiB packets. In contrast, ECI offers dramatically better latency even at 8 KiB packets.
Figure 10 shows transmit latency. In this case, PCIe PIO competes well with ECI across all packet sizes due to combining posted writes as we discussed in Section 4. PIO over ECI shows a slightly higher latency for small packets, due to it needing 2 round-trips over ECI for each 128 bytes, versus a single PCIe round trip. Both PIO solutions perform significantly better than DMA across all packet sizes.
Table 1 shows the tail latency impact for selected packet sizes for each of the NIC implementations. We select three representative packet sizes to evaluate tail latency: 64 for the smallest possible packet, 1536 for the normal Ethernet MTU, and 9600 for a common jumbo frame MTU. We use the tickless kernel setup mentioned earlier to avoid the 250 Hz Linux timer, in order to reveal the actual tail latency impact from our NIC.
We see that PIO over PCIe results in low tail latency compared to the descriptor-based approach, while ECI completely eliminates tail latency in this scenario.
These results show a clear advantage of using PIO over DMA, even for larger packets, and especially when using a coherent interconnect . The poor receive performance of PCIe-based PIO for receiving packets is due to the unposted-read limitation of PCIe, and disappears for transmit.
6.4. Accelerating Timely Dataflow
In this example we apply the techniques and results outlined so far to a practical workload: Timely Dataflow (Murray et al., 2013), a real-time stream-processing system. We compare the baseline performance to a partially FPGA-offloaded implementation under three CPU–FPGA communication scenarios: DMA over PCIe, PIO over PCIe, and PIO over ECI.
To test the limits, we construct a worst-case scenario: a large (31-element) graph of simple operators (filters). The large graph maximizes communication overhead in the form of progress tracking data, while the simple operators leave the FPGA with only a small performance lead over the CPU. We want to predict whether PIO can make FPGA offload work in regimes where the fixed overheads of DMA make it impractical, for example where small batch sizes are needed to ensure data freshness (as in Differential Dataflow (McSherry et al., 2013)).
@@ 6.4. Accelerating Timely Dataflow
As shown by the $r^{2}$ values, the fit for each dataset is excellent. We can thus model the time taken to process a batch as a fixed component, or overhead, independent of the amount of data processed (e.g., running the Timely scheduler, or setting up DMA descriptors) plus a component which scales linearly with the size of the batch.
@@
Additionally, bringing the fixed cost down reduces the smallest practical batch. If the relative overhead could be brought below 500ns/batch then with parallel cache lines giving a 1ns/B transfer cost, FPGA offload using ECI-PIO would outperform the CPU from a batch size of 128B. With such an optimized protocol, ECI-PIO would outperform all other cases until a batch size of 142KiB, above which PCIe-DMA would overtake it.
Figure 11. Timely Dataflow offload, graph size 31
@@
6.5. Offloading Bloom Filters
In a second acceleration experiment we offloaded Bloom filters (Bloom, 1970) to the FPGA. Bloom filters are used to efficiently check whether an element is a member of a set. It uses little memory and allows for fast lookups. A Bloom filter is equipped with $k$ different hash functions and a lookup table. Testing the presence of an element requires computing these $k$ hash functions and querying the lookup table with the results.
We took inspiration from the applications discussed in Fleet (Thomas et al., 2020). As the CPU baseline, we re-implemented the Fleet C++ code in Rust as a Timely Dataflow operator. In the Fleet implementation, the hashes were computed in parallel for each byte sequentially using AVX instructions, which we replicated in Rust using SIMD instructions. On the FPGA side, the hash computation is parallelized and the iteration over the element’s bytes is pipelined. We present the numbers of using a single thread on CPU and a single compute unit on FPGA in order to keep the comparison fair.
With the Bloom filter integrated into Timely Dataflow as a single offloaded operator, most of the additional overhead from processing sub-graphs is eliminated and the flow of data between FPGA and CPU is reduced to the minimal practical batch. In addition, the computational load is significantly higher compared to a simple filter. The $k$ hash functions consist of sequential bit shifting, addition, and XOR-ing, which are very fast in FPGA logic and deliver the offloading benefit comes from. Therefore in this application we concentrate on minimal practical batches and direct offloading benefits.
We implement the Bloom filter on the FPGA to take 128-byte elements with $k=8$ hash functions that take byte-wide inputs. We unroll the hash function calculations for each byte lane by a factor of 2, resulting in a 64-cycle latency for each element. We then pipeline the calculations with an initiation interval of 2 cycles to saturate the 512-bit bus. The return value is the set of 8 64-bit hashes for each element.
Instead of directly returning the hash results, the FPGA could have computed the lookup table positions or performed the lookup itself, returning either $k$ indices or binary values. In the case of indices, the read-to-write ratio would stay the same as in our approach, but for the lookup, the read requests would be minimized and writes would dominate. In this experiment, we assume that the lookup table is located elsewhere and focus on efficiently calculating the hash function and transferring the results to the lookup table.
In Figure 12, the latencies for PIO and CPU both start around 25 $\mu$s, due to the high overhead of streaming the input data. The real processing time of one element is 2.6 $\mu$s on the CPU and 1.7 $\mu$s when offloading with ECI. As overall processing time increases with the data size, we see the gap between CPU and PIO more clearly. After 8 KiB, same as in previous experiments, XDMA starts to perform better than other approaches as performance becomes dominated by the higher bulk throughput of PCIe DMA.
Figure 12. Offloading Bloom filters to the FPGA
@@ 7. Related work
Cho et al. (Cho et al., 2018) provide an excellent survey of existing CPU-device communication mechanisms, and address high-speed storage devices with microsecond latencies. They question the conventional wisdom that hiding the latency of interacting with such devices is not possible using existing techniques, and show how access latency using existing descriptor-based protocols can be effectively hidden by clever use of prefetching, better hardware queues, and user-level context switching. Our work is different (and complementary) in reducing the actual latency of each message between CPU registers and device memory.
7.1. Improving descriptor-based DMA
Cohort (Wei et al., 2023b) proposes a single, uniform, single-producer, single-consumer, queue-based DMA descriptor interface to all accelerators on an System-on-Chip (SoC) which replaces the variety of ad-hoc interfaces found on most SoCs. Cohort mandates an MMU on each accelerator, which maintains consistency with CPU MMUs and allows seamless user-space access to accelerator queues, and integrates with the cache coherence protocol to optimize the use of lock-free descriptor queues. We focus instead on what a latency-optimized interface might look like, eschewing descriptors in favor of direct cache line transfers.
Ensō (Sadok et al., 2023) shows the benefits of replacing the packet-based DMA software interface used by modern NICs with one based on a contiguous large buffer of opaque, variable-sized messages, with metadata sent via a different completion buffer. Ensō replaces descriptors with this buffer using an ingenious combination of direct MMIO writes to the PCIe device and polling memory queues written by the NIC. Our work similarly tries to reduce the overhead of descriptor management, but by focusing solely on latency-sensitive small payloads and avoiding DMA access completely.